Jar A contains one white and two black balls, jar B contains five white and two black balls, and jar C contains four white and two black balls. One ball is selected from each jar. What is the probability that the ball chosen from jar B will be white, given that exactly two black balls are selected?


Show your reasoning and calculations.


1
Expert's answer
2022-03-31T17:06:38-0400

A: 1w, 2b - 1

B: 5w, 2b - 1

C: 4w, 2b - 1


"P=\frac{2}{3}\frac{2}{6}\frac{5}{7}=\frac{10}{63}"